Help with quadratic equation with roots : √2x2 — 3x + √2 =0
Answer by Theo(13342)

if you graph this equation, you will get roots of:

x = sqrt(2)

x = sqrt(2)/2

sqrt(2) = 1.4142....

sqrt(2)/2 = .7071...

the graph is shown below:

###

if you use the quadratic formula, you will get the same answer.

my worksheet is shown below:

$$$

since the graph says those are the roots and the quadratic equations says those are the roots, i'm reasonably confident that those are the roots.

where the whole thing falls down is when i try to create factors from the roots.

then i run into difficulty.

but, since you were asked to find the roots, and not the factors, then i believe i've succeeded.

the roots are as indicated above as confirmed in the worksheet and the graph.